# Neural Acoustic Multipole Splatting for Room Impulse Response Synthesis

### v1

@@ 4.2 Ablation studies and discussions
We conducted ablation studies to investigate the efficacy of multipole splatting, by comparing the monopole and multipole models. The monopole model is implemented by setting the maximum order ($N$) of spherical harmonics to zero. We also considered two different pole initializations for each configuration. The dense initialization denotes the initialization described in Section 3.2, while the sparse initialization indicates the case where the number of points on each sphere is reduced such that the total number of poles is similar to that obtained after pruning the dense initialization model. The results are shown in Table 2.
For the same number of poles, multipole models consistently outperform monopole models across all datasets. Despite this, multipole models incur higher inference time due to their larger number of trainable parameters. Nevertheless, in both the Apartment 566 and 716 datasets, the sparse multipole model outperforms all monopole models, including the dense initialization model. This result indicates that multipoles offer a more expressive representation in complex acoustic environments. Conversely, in the MeshRIR dataset, both the multipole model with sparse initialization and the monopole model with dense initialization perform similarly, suggesting that monopoles might suffice for modeling uncomplicated, unobstructed rooms.
As indicated in Table 2, pruning decreased the multipoles to 225, 276, and 240 in the MeshRIR, Apartment 566, and Apartment 716 datasets, respectively, yet still outperformed the dense monopole model with 1089 poles. This highlights that the multipole model more effectively represents sound fields using only 20–22% of poles. The pruned model also resulted in over twice the inference speed compared to the non-pruned dense multipole initialization for each dataset. Moreover, the compact model aids in interpreting the physical structure of a sound field. To illustrate, we present the multipole distributions optimized for the MeshRIR dataset under three scenarios (Fig. 5): dense and sparse splatting without pruning, and dense splatting with pruning. In the pruned dense setup, two high-energy multipoles are placed near the source, whereas both dense and sparse setups without pruning spread many low-energy multipoles in the space. Thus, the pruning model offers a more interpretable, structured portrayal of a sound field using a limited set of multipoles with adaptable directivities and signal emission.
@@ 5 Conclusion
We introduced NAMS to synthesize room impulse responses via multipoles, using pruning to auto-select the optimal quantity. This approach outperforms existing methods while ensuring quick inference by efficiently representing sound fields with refined multipole directivities and placements. We demonstrated that multipoles offer a richer representation than monopoles in complex acoustic scenarios. Additionally, starting with a dense multipole set and optimizing through pruning achieves better results than manual initialization. This evidence shows NAMS efficiently represents RIRs. The next step for practical RIR estimation will be designing a generalized model for various source positions using fewer RIRs.

### v2

@@ 4.2 Ablation studies and discussions
We conducted ablation studies to investigate the efficacy of multipole splatting by comparing the monopole and multipole models. The monopole model is implemented by setting the maximum order ($N$) of spherical harmonics to zero. We also considered two different pole initializations for each configuration. The dense initialization denotes the initialization described in Section 3.2, while the sparse initialization indicates the case where the number of points on each sphere is reduced such that the total number of poles is similar to that obtained after pruning the dense initialization model. The results are shown in Table 2.
For the same number of poles, multipole models consistently outperform monopole models across all datasets. Despite this, multipole models incur longer inference time due to their larger number of trainable parameters. Nevertheless, in both the Apartment 566 and 716 datasets, the sparse multipole model outperforms all monopole models, including the dense initialization model. This result indicates that multipoles offer a more expressive representation in complex acoustic environments. Conversely, on the MeshRIR dataset, both the multipole model with sparse initialization and the monopole model with dense initialization perform similarly, suggesting that monopoles might suffice for modeling uncomplicated, unobstructed rooms.
As indicated in Table 2, pruning decreases the multipoles to 225, 276, and 240 in the MeshRIR, Apartment 566, and Apartment 716 datasets, respectively, while still outperforming the dense monopole model with 1,089 poles. This highlights that the multipole model more effectively represents sound fields using only 20–22% of poles. The pruned model also resulted in over twice the inference speed compared to the non-pruned dense multipole initialization for each dataset. Moreover, the compact model aids in interpreting the physical structure of a sound field. To illustrate, we present the multipole distributions optimized for the MeshRIR dataset under three scenarios (Fig. 5): dense and sparse splatting without pruning, and dense splatting with pruning. In the pruned dense setup, three high-energy multipoles are placed near the source, whereas both dense and sparse setups without pruning spread many low-energy multipoles in the space. Thus, the pruned model offers a more interpretable, structured portrayal of a sound field using a limited set of multipoles with adaptable directivities and signal emission. Additional demos are available at https://bgw6287.github.io/nams-demo/.
@@ 5 Conclusion
We introduced NAMS to synthesize room impulse responses via multipoles, using pruning to auto-select the optimal quantity. This approach outperforms existing methods while ensuring fast inference by efficiently representing sound fields with refined multipole directivities and placements. We demonstrated that multipoles offer a richer representation than monopoles in complex acoustic environments. Additionally, starting with a dense multipole set and optimizing through pruning achieves better results than manual initialization. This evidence shows NAMS efficiently represents RIRs. The next step for practical RIR estimation will be designing a generalized model for various source positions using fewer RIRs.